Question 2: You are given the following probability distribution for a stock: Probability Outcome .5 -6% 5 18% 1. A) Compute could you work out step by step how you got this answer ?
Answer #1

Expected return=Respective return*Respective probability

=(0.5*-6)+(0.5*18)=6%

probability Return probability*(return-Expected return)^2
0.5 -6 0.5*(-6-6)^2=72
0.5 18 0.5*(18-6)^2=72
Total=144%

Standard deviation=[Total probability*(return-Expected return)^2/Total probability]^(1/2)

=(144)^(1/2)

=12%

Coefficient of variation=Standard deviation/Expected return

=12/6=2